# Backpropagation-Free Test-Time Adaptation via Probabilistic Gaussian Alignment

### v5

@@ 1 Introduction
While Vision-language models (VLMs) such as CLIP CLIP can recognize diverse visual concepts using only class names, their robustness drops significantly when test-time distributions differ from those seen during pre-training TPT_NIPS22 ; C-TPT_ICLR24 . Test-time adaptation (TTA) offers a promising solution to enhance the robustness of VLMs by adapting them using only unlabeled test data wangtent . From a distributional perspective, TTA helps correct the mismatch between the class-conditional feature distributions learned during pretraining and those encountered at test time. Early approaches, including prompt tuning TPT_NIPS22 ; HisTPT_NIPS24 ; swapprompt_NIPS23 ; Promptalign_NIPS23 and adapter tuning methods Clip-adapter ; ape_iccv23 ; TaskRes_CVPR23 ; TPS_WACV25 ; DPE_NIPS24 , refine additional prompts or adapter layers through iterative optimization. While more parameter-efficient than full fine-tuning, these methods still depend on costly backpropagation, hindering their deployment in real-time or streaming scenarios Tip-adapter .
@@
Recent advances in TTA have increasingly emphasized efficiency, leading to the development of training-free approaches. Several methods leverage memory banks constructed from test samples TDA_CVPR24 ; Boostadapter_NIPS24 or few-shot training data to adapt classifiers based on sample similarities Tip-adapter ; DMN_CVPR24 , but often require tuning task-specific fusion coefficients, limiting their generalizability across tasks. Other methods Zlap_CVPR24 ; ECALP_ICLR25 propagate labels from text prompts to test samples via static graph construction, yet frequently depend on external datasets and involve costly iterative updates.
@@
Alongside these trends, transductive learning transdu_NIPS03 has also gained attention. This setting assumes full or partial access to the target data during inference, enabling the model to capture the global structure of the test distribution. It works well in offline scenarios such as batch processing of medical scans or document classification, where the model can benefit from support sets or feature clusters. However, such methods often require complete access to target data TransCLIP_NIPS24 ; Frolic_NIPS24 ; Histo_TransCLIP ; StatA_CVPR25 , or even source data hard_ICLR24 ; Zlap_CVPR24 , making them unsuitable for strict online or streaming environments where samples arrive and are processed sequentially, one by one, without access to future data.
@@
Despite advances in efficiency and data accessibility, existing TTA approaches still overlook how class-conditional feature distributions are structured. Many rely solely on text-based prototypes to define decision boundaries TPT_NIPS22 ; TPS_WACV25 ; DPE_NIPS24 ; Clip-adapter ; Promptalign_NIPS23 or adjust similarity scores using task-specific scaling TDA_CVPR24 ; Boostadapter_NIPS24 ; DMN_CVPR24 ; Tip-adapter ; Sus-x_ICCV23 , without explicit class distribution modeling. Moreover, since TTA operates without supervision on the target domain and lacks access to source data, estimating class distributions remains challenging. As a result, existing methods often fail to capture intra-class variation and inter-class confusion (see Figure 3), which leads to unstable decision boundaries and reduced robustness under distribution shifts.
@@
Prompt Tuning TPT_NIPS22 ; HisTPT_NIPS24 ; swapprompt_NIPS23 ; Promptalign_NIPS23 ; dynaprompt_ICLR25
Adapter Tuning Clip-adapter ; ape_iccv23 ; TaskRes_CVPR23 ; TPS_WACV25 ; DPE_NIPS24
Similarity Score TDA_CVPR24 ; Boostadapter_NIPS24 ; DMN_CVPR24 ; Tip-adapter ; Sus-x_ICCV23
Transductive Learning Zlap_CVPR24 ; ECALP_ICLR25 ; TransCLIP_NIPS24 ; Frolic_NIPS24 ; hard_ICLR24
@@ 2 Related Works
Online TTA adapts pre-trained VLMs to downstream tasks in streaming scenarios. Existing approaches fall into two main directions: prompt tuning and adapter tuning. Prompt-based methods TPT_NIPS22 ; swapprompt_NIPS23 ; Promptalign_NIPS23 ; SelfTPT ; HisTPT_NIPS24 ; DiffTPT_ICCV23 ; C-TPT_ICLR24 adapt prompts by minimizing entropy across augmented views TPT_NIPS22 ; DiffTPT_ICCV23 ; swapprompt_NIPS23 or aligning multi-modal features using proxy data Promptalign_NIPS23 . Adapter-based approaches Clip-adapter ; DPE_NIPS24 ; TPS_WACV25 inject lightweight modules into VLMs while keeping the backbone frozen. Despite their effectiveness, these methods rely on backpropagation, incurring high computation and memory costs. Recently, training-free TTA methods have gained attention for their efficiency. Approaches like Tip-Adapter Tip-adapter and SuS-X Sus-x_ICCV23 utilize class-wise support samples from external sources for nearest-neighbor matching. Others TDA_CVPR24 ; DMN_CVPR24 ; Boostadapter_NIPS24 adapt predictions by retrieving historical features based on sample similarity. While gradient-free, they often require careful hyperparameter tuning and lack explicit modeling of class distributions, making them sensitive to distribution shifts and prone to unstable decision boundaries. We propose a distribution-aware and propagation-free TTA paradigm, achieving stable and efficient zero-shot performance in challenging scenarios involving severe distribution shifts and fine-grained tasks.
@@
Transductive learning transductive_icml99 ; transdu_NIPS03 , where all test samples are accessible during inference, was initially explored in the few-shot learning literature for leveraging both labeled and unlabeled samples to improve generalization. By modeling the structure of the test set, it often outperforms inductive methods transduc_fs ; laplacian ; TIM . Inspired by these successes, recent efforts have extended transductive learning to the multimodal domain to improve the zero-shot adaptation of VLMs. ZLaP Zlap_CVPR24 and ECALP ECALP_ICLR25 build similarity graphs over test-time features to propagate labels but rely on costly iterative refinement steps. Wang et al. hard_ICLR24 employs a Gaussian Discriminant Analysis classifier by estimating distributional statistics from external source data. Frolic Frolic_NIPS24 adjusts predictions by learning class-wise prompt distributions inferred from target data moments. TransCLIP TransCLIP_NIPS24 , Histo-TransCLIP Histo_TransCLIP and StatA StatA_CVPR25 model each class with a multivariate Gaussian and penalize the deviations from pseudo-labels and text-driven priors. While effective, these methods often require access to all or large batches of test data, limiting their use in online or streaming settings. To overcome this, we propose an adaptive method that builds class-wise Gaussian distributions using reliably accumulated high-confidence predictions. This supports both transductive and online adaptation, offering strong deployment flexibility.
@@ 3.2 ADAPT: Backpropagation-free and Distribution-aware Test-time Adaptation
Gaussian Discriminant Analysis (GDA) is a classical generative model that assigns class labels based on the likelihood of class-conditional distributions GDA_96 . Recent studies have shown that CLIP features exhibit class-conditional clustering and can be well approximated by Gaussian distributions hard_ICLR24 ; Frolic_NIPS24 ; TransCLIP_NIPS24 . Motivated by these findings, we adopt GDA as a backpropagation-free probabilistic classifier for test-time adaptation (TTA). Specifically, to achieve efficient, real-time adaptation, we simply assume that features conditioned on class $k$ follow a Gaussian distribution with a shared covariance matrix:
@@
While GDA enables label prediction based on estimated class statistics, directly applying it in an online test-time setting results in biased likelihood estimates due to limited samples and overconfident early predictions. Inspired by prior works on transductive adaptation and regularized inference laplacian ; TransCLIP_NIPS24 ; StatA_CVPR25 , we propose a bias correction framework that minimizes a regularized objective combining three components: (i) Online Negative Log-Likelihood $z_{i}^{\top}\log\mathbb{P}_{i}$, (ii) CLIP Prior-based Regularization $\mathcal{R}(z_{i};\hat{y}_{i})$, and (iii) Knowledge Bank-guided Consistency Regularization $\mathcal{R}(z_{i};\mathcal{B})$, as follows:
@@
In high-dimensional feature spaces, directly inverting the empirical covariance matrix is often ill-conditioned and leads to biased estimates hard_ICLR24 . To address this, we adopt a Bayesian ridge estimator Bayes_ridge_estimator with shrinkage regularization to approximate the shared covariance:
@@ 4 Experiments
Online CLIP CLIP - 66.74 47.79 60.89 73.99 46.12 57.20 59.11
Tip-Adapter Tip-adapter 70.75 51.04 63.41 77.76 48.88 60.27 62.37
TPT TPT_NIPS22 68.98 54.77 63.45 77.06 47.97 60.81 62.45
DiffTPT DiffTPT_ICCV23 70.30 55.68 65.10 75.00 46.80 60.65 62.58
C-TPT C-TPT_ICLR24 68.50 51.60 62.70 76.00 47.90 59.55 61.34
DMN DMN_CVPR24 72.25 58.28 65.17 78.55 53.20 63.80 65.49
DPE DPE_NIPS24 71.91 59.63 65.44 80.40 52.26 64.43 65.93
TPS TPS_WACV25 70.38 59.21 63.80 77.49 49.57 62.52 64.09
DynaPrompt dynaprompt_ICLR25 69.61 56.17 64.67 78.17 48.22 61.81 63.37
B2TPT B2tpt 69.57 55.26 65.40 78.64 49.53 62.21 63.68
MTA MTA_CVPR24 70.08 58.06 64.24 78.33 49.61 62.56 64.06
TDA TDA_CVPR24 69.51 60.11 64.67 80.24 50.54 63.89 65.01
ZERO zero_NIPS24 69.31 59.61 64.16 77.22 48.40 62.35 63.74
AWT AWT_NIPS24 71.32 60.33 65.15 80.64 51.60 64.43 65.81
RA-TTA RA-TTA__ICLR25 70.58 59.21 64.16 79.68 50.83 63.47 64.89
BCA BCA_CVPR25 70.22 61.14 64.90 80.72 50.87 64.41 65.57
TCA TCA_arxiv25 68.88 50.13 62.10 77.11 48.95 59.57 61.43
Dota dota_arXiv24 70.68 61.19 64.41 81.17 51.33 64.53 65.76
@@
Trans. GDA-CLIP hard_ICLR24 64.13 19.72 55.67 55.30 34.32 41.25 45.83
TransCLIP TransCLIP_NIPS24 70.30 49.50 62.30 75.00 49.70 59.13 61.36
Frolic Frolic_NIPS24 70.90 60.40 64.70 80.70 53.30 64.78 66.00
TIMO TIMO_AAAI25 64.63 22.06 56.40 58.47 35.96 43.22 47.50
@@
Online CLIP CLIP 24.25 15.71 24.46 22.60 33.08 31.06 37.61 55.62 17.11 13.43 33.04 33.70 13.25 14.16 13.48 25.50
TPT TPT_NIPS22 27.56 15.48 26.16 26.94 36.74 34.28 39.38 60.22 16.96 15.64 40.74 37.90 10.64 11.94 10.92 27.43
DiffTPT DiffTPT_ICCV23 25.63 16.96 26.74 25.40 35.99 34.57 39.83 59.01 17.32 17.16 38.43 35.47 12.97 13.60 13.21 27.49
TDA TDA_CVPR24 26.53 17.91 27.35 25.90 36.50 34.84 40.53 58.57 20.16 16.62 35.65 36.69 15.42 16.46 16.03 28.34
DMN DMN_CVPR24 26.06 17.19 26.61 25.23 34.81 33.48 38.93 58.70 19.38 15.40 35.32 36.49 14.33 15.33 14.69 27.46
@@
Trans. ZLaP Zlap_CVPR24 24.88 16.13 25.77 24.36 34.43 32.63 38.56 58.42 17.53 14.21 33.72 35.52 12.83 14.03 13.27 26.42
TransCLIP TransCLIP_NIPS24 25.35 16.40 25.53 23.22 34.58 32.47 39.65 59.04 17.72 14.76 35.22 35.53 14.82 16.11 15.60 27.07
StatA StatA_CVPR25 20.23 13.29 20.38 18.84 31.30 29.80 34.58 54.79 11.24 11.80 26.31 33.20 9.58 10.52 10.12 22.40
@@ 4.1 Setup
We evaluated the proposed ADAPT on three different tasks: natural distribution shift, fine-grained categorization, and corruption robustness. Specifically, for natural distribution shift, we use multiple datasets including ImageNet imagenet , ImageNet-A imagenet-a , ImageNet-R imagenet-r ,ImageNet-V imagenet-v , and ImageNet-Sketch imagenet-sketch . The corruption robustness task is evaluted on ImageNet-C imagenet-c , which contains 15 corruption types covering noise, blur, weather, and digital artifacts. We also evaluate performance on 10 fine-grained recognition datasets: Aircraft aircraft , Caltech101 caltech101 , Cars cars , DTD dtd , EuroSAT eurosat , Flower102 flowers102 , Food101 food101 , Pets pets , SUN397 sun397 and UCF101 ucf101 .
@@
We evaluate ADAPT under two TTA protocols: Online (batch size=1) and Transductive. In the online setting, the model processes one sample at a time with access to current and past inputs, following the sequential setup in TPT_NIPS22 ; TDA_CVPR24 . In the transductive setting TransCLIP_NIPS24 ; Histo_TransCLIP , all test samples are accessible at once. We use CLIP CLIP ViT-B/16 as the backbone. We set the coefficient $\alpha$ to 0.9, and assign the knowledge bank size $L$ as 16 for online and 6 for transductive evaluation. All experiments are conducted on an NVIDIA RTX A6000 GPU.
@@
Online CLIP CLIP - 23.70 92.98 65.24 44.44 41.42 67.28 83.80 87.98 62.55 65.08 63.45
TPT TPT_NIPS22 24.78 94.16 66.87 47.75 42.44 68.98 84.67 87.79 65.50 68.04 65.10
DiffTPT DiffTPT_ICCV23 25.60 92.49 67.01 47.00 43.13 70.10 87.23 88.22 65.74 68.22 65.47
C-TPT C-TPT_ICLR24 24.00 93.60 65.80 46.00 43.20 79.80 83.70 88.20 64.80 65.70 64.48
DMN DMN_CVPR24 30.03 95.38 67.96 55.85 59.43 74.49 85.08 92.04 70.18 72.51 70.30
TPS ECALP_ICLR25 26.27 94.56 67.00 53.80 42.11 71.69 84.78 87.82 68.25 71.18 66.75
DPE DPE_NIPS24 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 69.40
HisTPT HisTPT_NIPS24 26.90 94.50 69.20 48.90 49.70 71.20 89.30 89.10 67.20 70.10 67.61
DynaPrompt dynaprompt_ICLR25 24.33 94.32 67.65 47.96 42.28 69.95 85.42 88.28 66.32 68.72 65.52
MTA MTA_CVPR24 25.32 94.13 66.36 45.59 38.71 68.26 84.95 88.22 64.98 68.11 64.46
TDA TDA_CVPR24 23.91 94.24 67.28 47.40 58.00 71.42 86.14 88.63 67.62 70.66 67.53
ZLaP Zlap_CVPR24 25.40 93.10 65.60 48.60 55.60 73.50 86.90 87.10 67.40 71.50 67.47
ZERO zero_NIPS24 25.21 93.66 68.04 46.12 34.33 67.68 86.53 87.75 65.03 67.77 64.21
BCA BCA_CVPR25 28.59 94.69 66.86 53.49 56.63 73.12 85.97 90.43 68.41 67.59 68.58
OGA OGA__arxiv25 23.20 93.60 68.10 47.90 54.20 69.20 85.60 89.40 67.90 71.40 67.05
TCA TCA_arxiv25 24.87 93.63 65.33 46.16 70.43 73.33 85.31 89.53 65.92 72.38 68.69
Dota dota_arXiv24 25.59 94.32 69.48 47.87 57.65 74.67 87.02 91.69 69.70 72.06 69.01
@@
Trans. GDA-CLIP hard_ICLR24 18.69 87.53 60.78 46.81 49.92 72.65 78.25 89.90 63.60 68.70 63.68
ZLaP Zlap_CVPR24 26.30 91.80 66.80 46.00 57.70 67.90 87.20 87.90 67.80 73.80 67.32
TransCLIP TransCLIP_NIPS24 26.90 92.70 69.40 49.50 65.10 76.70 87.10 92.60 68.90 74.40 70.33
Frolic Frolic_NIPS24 31.40 95.10 69.10 56.10 58.50 74.80 87.10 92.90 70.80 75.20 71.10
StatA StatA_CVPR25 24.70 94.20 68.00 48.40 67.30 75.20 87.10 92.40 68.70 73.50 69.95

### v8

@@ 1 Introduction
While Vision-language models (VLMs) such as CLIP Radford et al. (2021) can recognize diverse visual concepts using only class names, their robustness drops significantly when test-time distributions differ from those seen during pre-training Manli et al. (2022); Yoon et al. (2024). Test-time adaptation (TTA) offers a promising solution to enhance the robustness of VLMs by adapting them using only unlabeled test data Wang et al. (2021). From a distributional perspective, TTA helps correct the mismatch between the class-conditional feature distributions learned during pretraining and those encountered at test time. Early approaches, including prompt tuning Manli et al. (2022); Zhang et al. (2024b); Ma et al. (2023); Samadh et al. (2023) and adapter tuning methods Gao et al. (2024); Zhu et al. (2023); Yu et al. (2023); Sui et al. (2025); Zhang et al. (2024a), refine additional prompts or adapter layers through iterative optimization. While more parameter-efficient than full fine-tuning, these methods still depend on costly backpropagation, hindering their deployment in real-time or streaming scenarios Zhang et al. (2022).
@@
Recent advances in TTA have increasingly emphasized efficiency, leading to the development of training-free approaches. Several methods leverage memory banks constructed from test samples Karmanov et al. (2024); Zhang et al. (2024c) or few-shot training data to adapt classifiers based on sample similarities Zhang et al. (2022, 2024d), but often require tuning task-specific fusion coefficients, limiting their generalizability across tasks. Other methods Kalantidis et al. (2024); Li et al. (2025b) propagate labels from text prompts to test samples via static graph construction, yet frequently depend on external datasets and involve costly iterative updates.
@@
Alongside these trends, transductive learning Zhou et al. (2003) has also gained attention. This setting assumes full or partial access to the target data during inference, enabling the model to capture the global structure of the test distribution. It works well in offline scenarios such as batch processing of medical scans or document classification, where the model can benefit from support sets or feature clusters. However, such methods often require complete access to target data Zanella et al. (2024a); Zhu et al. (2024a); Zanella et al. (2024b, 2025), or even source data Wang et al. (2024); Kalantidis et al. (2024), making them unsuitable for strict online or streaming environments where samples arrive and are processed sequentially, one by one, without access to future data.
@@
Despite advances in efficiency and data accessibility, existing TTA approaches still overlook how class-conditional feature distributions are structured. Many rely solely on text-based prototypes to define decision boundaries Manli et al. (2022); Sui et al. (2025); Zhang et al. (2024a); Gao et al. (2024); Samadh et al. (2023) or adjust similarity scores using task-specific scaling Karmanov et al. (2024); Zhang et al. (2024c, d, 2022); Udandarao et al. (2023), without explicit class distribution modeling. Moreover, since TTA operates without supervision on the target domain and lacks access to source data, estimating class distributions remains challenging. As a result, existing methods often fail to capture intra-class variation and inter-class confusion (see Figure 3), which leads to unstable decision boundaries and reduced robustness under distribution shifts.
@@
Prompt Tuning Manli et al. (2022); Zhang et al. (2024b); Ma et al. (2023); Samadh et al. (2023); Xiao et al. (2025)
Adapter Tuning Gao et al. (2024); Zhu et al. (2023); Yu et al. (2023); Sui et al. (2025); Zhang et al. (2024a)
Similarity Score Karmanov et al. (2024); Zhang et al. (2024c, d, 2022); Udandarao et al. (2023)
Transductive Learning Kalantidis et al. (2024); Li et al. (2025b); Zanella et al. (2024a); Zhu et al. (2024a); Wang et al. (2024)
@@ 2 Related Works
Online TTA adapts pre-trained VLMs to downstream tasks in streaming scenarios. Existing approaches fall into two main directions: prompt tuning and adapter tuning. Prompt-based methods Manli et al. (2022); Ma et al. (2023); Samadh et al. (2023); Zhu et al. (2024c); Zhang et al. (2024b); Feng et al. (2023); Yoon et al. (2024) adapt prompts by minimizing entropy across augmented views Manli et al. (2022); Feng et al. (2023); Ma et al. (2023) or aligning multi-modal features using proxy data Samadh et al. (2023). Adapter-based approaches Gao et al. (2024); Zhang et al. (2024a); Sui et al. (2025) inject lightweight modules into VLMs while keeping the backbone frozen. Despite their effectiveness, these methods rely on backpropagation, incurring high computation and memory costs. Recently, training-free TTA methods have gained attention for their efficiency. Approaches like Tip-Adapter Zhang et al. (2022) and SuS-X Udandarao et al. (2023) utilize class-wise support samples from external sources for nearest-neighbor matching. Others Karmanov et al. (2024); Zhang et al. (2024d, c) adapt predictions by retrieving historical features based on sample similarity. While gradient-free, they often require careful hyperparameter tuning and lack explicit modeling of class distributions, making them sensitive to distribution shifts and prone to unstable decision boundaries. We propose a distribution-aware and propagation-free TTA paradigm, achieving stable and efficient zero-shot performance in challenging scenarios involving severe distribution shifts and fine-grained tasks.
@@
Transductive learning Joachims (1999); Zhou et al. (2003), where all test samples are accessible during inference, was initially explored in the few-shot learning literature for leveraging both labeled and unlabeled samples to improve generalization. By modeling the structure of the test set, it often outperforms inductive methods Liu et al. (2019); Ziko et al. (2020); Boudiaf et al. (2020). Inspired by these successes, recent efforts have extended transductive learning to the multimodal domain to improve the zero-shot adaptation of VLMs. ZLaP Kalantidis et al. (2024) and ECALP Li et al. (2025b) build similarity graphs over test-time features to propagate labels but rely on costly iterative refinement steps. Wang et al. Wang et al. (2024) employs a Gaussian Discriminant Analysis classifier by estimating distributional statistics from external source data. Frolic Zhu et al. (2024a) adjusts predictions by learning class-wise prompt distributions inferred from target data moments. TransCLIP Zanella et al. (2024a), Histo-TransCLIP Zanella et al. (2024b) and StatA Zanella et al. (2025) model each class with a multivariate Gaussian and penalize the deviations from pseudo-labels and text-driven priors. While effective, these methods often require access to all or large batches of test data, limiting their use in online or streaming settings. To overcome this, we propose an adaptive method that builds class-wise Gaussian distributions using reliably accumulated high-confidence predictions. This supports both transductive and online adaptation, offering strong deployment flexibility.
@@ 3.2 ADAPT: Backpropagation-free and Distribution-aware Test-time Adaptation
Gaussian Discriminant Analysis (GDA) is a classical generative model that assigns class labels based on the likelihood of class-conditional distributions Hastie and Tibshirani (1996). Recent studies have shown that CLIP features exhibit class-conditional clustering and can be well approximated by Gaussian distributions Wang et al. (2024); Zhu et al. (2024a); Zanella et al. (2024a). Motivated by these findings, we adopt GDA as a backpropagation-free probabilistic classifier for test-time adaptation (TTA). Specifically, to achieve efficient, real-time adaptation, we simply assume that features conditioned on class $k$ follow a Gaussian distribution with a shared covariance matrix:
@@
While GDA enables label prediction based on estimated class statistics, directly applying it in an online test-time setting results in biased likelihood estimates due to limited samples and overconfident early predictions. Inspired by prior works on transductive adaptation and regularized inference Ziko et al. (2020); Zanella et al. (2024a, 2025), we propose a bias correction framework that minimizes a regularized objective combining three components: (i) Online Negative Log-Likelihood $z_{i}^{\top}\log\mathbb{P}_{i}$, (ii) CLIP Prior-based Regularization $\mathcal{R}(z_{i};\hat{y}_{i})$, and (iii) Knowledge Bank-guided Consistency Regularization $\mathcal{R}(z_{i};\mathcal{B})$, as follows:
@@
In high-dimensional feature spaces, directly inverting the empirical covariance matrix is often ill-conditioned and leads to biased estimates Wang et al. (2024). To address this, we adopt a Bayesian ridge estimator Kubokawa and Srivastava (2008) with shrinkage regularization to approximate the shared covariance:
@@ 4 Experiments
Online CLIP Radford et al. (2021) - 66.74 47.79 60.89 73.99 46.12 57.20 59.11
Tip-Adapter Zhang et al. (2022) 70.75 51.04 63.41 77.76 48.88 60.27 62.37
TPT Manli et al. (2022) 68.98 54.77 63.45 77.06 47.97 60.81 62.45
DiffTPT Feng et al. (2023) 70.30 55.68 65.10 75.00 46.80 60.65 62.58
C-TPT Yoon et al. (2024) 68.50 51.60 62.70 76.00 47.90 59.55 61.34
DMN Zhang et al. (2024d) 72.25 58.28 65.17 78.55 53.20 63.80 65.49
DPE Zhang et al. (2024a) 71.91 59.63 65.44 80.40 52.26 64.43 65.93
TPS Sui et al. (2025) 70.38 59.21 63.80 77.49 49.57 62.52 64.09
DynaPrompt Xiao et al. (2025) 69.61 56.17 64.67 78.17 48.22 61.81 63.37
B2TPT Meng et al. (2025) 69.57 55.26 65.40 78.64 49.53 62.21 63.68
MTA Zanella and Ben Ayed (2024) 70.08 58.06 64.24 78.33 49.61 62.56 64.06
TDA Karmanov et al. (2024) 69.51 60.11 64.67 80.24 50.54 63.89 65.01
ZERO Farina et al. (2024) 69.31 59.61 64.16 77.22 48.40 62.35 63.74
AWT Zhu et al. (2024b) 71.32 60.33 65.15 80.64 51.60 64.43 65.81
RA-TTA Lee et al. (2025) 70.58 59.21 64.16 79.68 50.83 63.47 64.89
BCA Zhou et al. (2025) 70.22 61.14 64.90 80.72 50.87 64.41 65.57
TCA Wang et al. (2025) 68.88 50.13 62.10 77.11 48.95 59.57 61.43
Dota Han et al. (2024) 70.68 61.19 64.41 81.17 51.33 64.53 65.76
@@
Trans. GDA-CLIP Wang et al. (2024) 64.13 19.72 55.67 55.30 34.32 41.25 45.83
TransCLIP Zanella et al. (2024a) 70.30 49.50 62.30 75.00 49.70 59.13 61.36
Frolic Zhu et al. (2024a) 70.90 60.40 64.70 80.70 53.30 64.78 66.00
TIMO Li et al. (2025a) 64.63 22.06 56.40 58.47 35.96 43.22 47.50
@@
Online CLIP Radford et al. (2021) 24.25 15.71 24.46 22.60 33.08 31.06 37.61 55.62 17.11 13.43 33.04 33.70 13.25 14.16 13.48 25.50
TPT Manli et al. (2022) 27.56 15.48 26.16 26.94 36.74 34.28 39.38 60.22 16.96 15.64 40.74 37.90 10.64 11.94 10.92 27.43
DiffTPT Feng et al. (2023) 25.63 16.96 26.74 25.40 35.99 34.57 39.83 59.01 17.32 17.16 38.43 35.47 12.97 13.60 13.21 27.49
TDA Karmanov et al. (2024) 26.53 17.91 27.35 25.90 36.50 34.84 40.53 58.57 20.16 16.62 35.65 36.69 15.42 16.46 16.03 28.34
DMN Zhang et al. (2024d) 26.06 17.19 26.61 25.23 34.81 33.48 38.93 58.70 19.38 15.40 35.32 36.49 14.33 15.33 14.69 27.46
@@
Trans. ZLaP Kalantidis et al. (2024) 24.88 16.13 25.77 24.36 34.43 32.63 38.56 58.42 17.53 14.21 33.72 35.52 12.83 14.03 13.27 26.42
TransCLIP Zanella et al. (2024a) 25.35 16.40 25.53 23.22 34.58 32.47 39.65 59.04 17.72 14.76 35.22 35.53 14.82 16.11 15.60 27.07
StatA Zanella et al. (2025) 20.23 13.29 20.38 18.84 31.30 29.80 34.58 54.79 11.24 11.80 26.31 33.20 9.58 10.52 10.12 22.40
@@ 4.1 Setup
We evaluated the proposed ADAPT on three different tasks: natural distribution shift, fine-grained categorization, and corruption robustness. Specifically, for natural distribution shift, we use multiple datasets including ImageNet Deng et al. (2009), ImageNet-A Hendrycks et al. (2021b), ImageNet-R Hendrycks et al. (2021a),ImageNet-V Recht et al. (2019), and ImageNet-Sketch Wang et al. (2019). The corruption robustness task is evaluted on ImageNet-C Hendrycks and Dietterich (2019), which contains 15 corruption types covering noise, blur, weather, and digital artifacts. We also evaluate performance on 10 fine-grained recognition datasets: Aircraft Maji et al. (2013), Caltech101 Fei-Fei et al. (2004), Cars Krause et al. (2013), DTD Cimpoi et al. (2014), EuroSAT Helber et al. (2019), Flower102 Nilsback and Zisserman (2008), Food101 Bossard et al. (2014), Pets Parkhi et al. (2012), SUN397 Xiao et al. (2010) and UCF101 Soomro et al. (2012).
@@
We evaluate ADAPT under two TTA protocols: Online (batch size=1) and Transductive. In the online setting, the model processes one sample at a time with access to current and past inputs, following the sequential setup in Manli et al. (2022); Karmanov et al. (2024). In the transductive setting Zanella et al. (2024a, b), all test samples are accessible at once. We use CLIP Radford et al. (2021) ViT-B/16 as the backbone. We set the coefficient $\alpha$ to 0.9, and assign the knowledge bank size $L$ as 16 for online and 6 for transductive evaluation. All experiments are conducted on an NVIDIA RTX A6000 GPU.
@@
Online CLIP Radford et al. (2021) - 23.70 92.98 65.24 44.44 41.42 67.28 83.80 87.98 62.55 65.08 63.45
TPT Manli et al. (2022) 24.78 94.16 66.87 47.75 42.44 68.98 84.67 87.79 65.50 68.04 65.10
DiffTPT Feng et al. (2023) 25.60 92.49 67.01 47.00 43.13 70.10 87.23 88.22 65.74 68.22 65.47
C-TPT Yoon et al. (2024) 24.00 93.60 65.80 46.00 43.20 79.80 83.70 88.20 64.80 65.70 64.48
DMN Zhang et al. (2024d) 30.03 95.38 67.96 55.85 59.43 74.49 85.08 92.04 70.18 72.51 70.30
TPS Li et al. (2025b) 26.27 94.56 67.00 53.80 42.11 71.69 84.78 87.82 68.25 71.18 66.75
DPE Zhang et al. (2024a) 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 69.40
HisTPT Zhang et al. (2024b) 26.90 94.50 69.20 48.90 49.70 71.20 89.30 89.10 67.20 70.10 67.61
DynaPrompt Xiao et al. (2025) 24.33 94.32 67.65 47.96 42.28 69.95 85.42 88.28 66.32 68.72 65.52
MTA Zanella and Ben Ayed (2024) 25.32 94.13 66.36 45.59 38.71 68.26 84.95 88.22 64.98 68.11 64.46
TDA Karmanov et al. (2024) 23.91 94.24 67.28 47.40 58.00 71.42 86.14 88.63 67.62 70.66 67.53
ZLaP Kalantidis et al. (2024) 25.40 93.10 65.60 48.60 55.60 73.50 86.90 87.10 67.40 71.50 67.47
ZERO Farina et al. (2024) 25.21 93.66 68.04 46.12 34.33 67.68 86.53 87.75 65.03 67.77 64.21
BCA Zhou et al. (2025) 28.59 94.69 66.86 53.49 56.63 73.12 85.97 90.43 68.41 67.59 68.58
OGA Fuchs et al. (2025) 23.20 93.60 68.10 47.90 54.20 69.20 85.60 89.40 67.90 71.40 67.05
TCA Wang et al. (2025) 24.87 93.63 65.33 46.16 70.43 73.33 85.31 89.53 65.92 72.38 68.69
Dota Han et al. (2024) 25.59 94.32 69.48 47.87 57.65 74.67 87.02 91.69 69.70 72.06 69.01
@@
Trans. GDA-CLIP Wang et al. (2024) 18.69 87.53 60.78 46.81 49.92 72.65 78.25 89.90 63.60 68.70 63.68
ZLaP Kalantidis et al. (2024) 26.30 91.80 66.80 46.00 57.70 67.90 87.20 87.90 67.80 73.80 67.32
TransCLIP Zanella et al. (2024a) 26.90 92.70 69.40 49.50 65.10 76.70 87.10 92.60 68.90 74.40 70.33
Frolic Zhu et al. (2024a) 31.40 95.10 69.10 56.10 58.50 74.80 87.10 92.90 70.80 75.20 71.10
StatA Zanella et al. (2025) 24.70 94.20 68.00 48.40 67.30 75.20 87.10 92.40 68.70 73.50 69.95